OpenStudy (anonymous):

A piece of art is in the shape of an equilateral triangle with sides of 6 in. Find the area of the piece of art. Round your answer to the nearest tenth.? A. 12.7 in.2 B. 15.6 in.2 C. 31.2 in.2 D. none of these

OpenStudy (anonymous):

A=.5(B*H). B=3, H=6. A=.5(6*3)=9 in 2

OpenStudy (anonymous):

A=.5(3*6)=9in^2 a=90

OpenStudy (anonymous):

hey sorry let me explain it to u

OpenStudy (anonymous):

so area of triangle is .5*base*height right

OpenStudy (anonymous):

we know our base, because we can divide 6 by 2 to get 3, but we dont know the height

OpenStudy (anonymous):

so u need to use pythgorean thm. A^2+B^2=C^2

OpenStudy (anonymous):

A being height. A^2+3^2=6^2

OpenStudy (anonymous):

Solve for A, you will get 5.2

OpenStudy (anonymous):

so A=.5*5.2*3=7.79

OpenStudy (anonymous):

but now multiply by two because that is just half of the triangle, so its 15.6 in2

OpenStudy (anonymous):

A is also area